# BroadGen: A Framework for Generating Effective and Efficient Advertiser Broad Match Keyphrase Recommendations

### v2

@@
Abstract.
In the domain of sponsored search advertising, the focus of Keyphrase recommendation has largely been on exact match types, which pose issues such as high management expenses, limited targeting scope, and evolving search query patterns. Alternatives like Broad match types can alleviate certain drawbacks of exact matches but present challenges like poor targeting accuracy and minimal supervisory signals owing to limited advertiser usage. This research defines the criteria for an ideal broad match, emphasizing on both efficiency and effectiveness, ensuring that a significant portion of matched queries are relevant. We propose BroadGen, an innovative framework that recommends efficient and effective broad match keyphrases by utilizing historical search query data. Additionally, we demonstrate that BroadGen, through token correspondence modeling, maintains better query stability over time. BroadGen’s capabilities allow it to serve daily, millions of sellers at eBay with over 2.3 billion items.
Information Retrieval, Sponsored search advertising
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 34th ACM Conference on Information and Knowledge Management ; Nov 10–14, 2025; Seoulisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systemsccs: Information systems Retrieval effectiveness
1. Introduction
Advertisers or sellers on e-commerce platforms bid on keyphrases recommended by the platform as well as their own custom keyphrases to increase their sales (Figure 1). Keyphrase recommendation is the subject of a lot of research and well studied in several works (Kharbanda et al., 2025; Mishra et al., 2025, 2024; Dahiya et al., 2021, 2023a; Jain et al., 2023; Dahiya et al., 2023b; You et al., 2019) due to their challenging research and industrial importance. This type of recommendation focuses only on Exact matches. The recommendation to the seller, is termed as Keyphrase and the buyer input search is termed as Query. The platform matches the (recommended) keywords bid by the sellers to the buyer search queries (Dey et al., 2025b).
@@ 1. Introduction
Figure 1. Screenshot of keyphrases for manual targeting in Promoted Listings Priority for eBay Advertising.
Each keyphrase is tagged with a Match Type; for this work we define them as:
  • Exact match type - The keyphrase is matched exactly to the query accommodating different cases, plurals, contractions, and similar substitutions.
  • Phrase match type - In addition to the criteria of exact match it is more flexible by allowing modification to the keyphrase by in-position tokens expansion but maintaining the token order.
@@
For example, if a query is ‘‘reebok men’s shoes size 9’’, then the exact matching keyphrases can be ‘‘reebok men shoes size 9’’ or ‘‘Reebok men’s shoe size 9’’. The phrase match can be ‘‘reebok men shoe’’ while broad match can be ‘‘reebok shoe men’’. More examples of how different match types keyphrases for an item are matched to queries are shown in Figure 2. A keyphrase with different match types can be recommended to sellers who can bid on it to increase their item visibility depending on buyer searches.
Figure 2. Matching architecture
1.1. Problems with Exact Match Type
1.1.1. Shifting label distributions
@@
Our analysis reveals that eBay’s search queries adhere to the 80/20 rule like other platforms (Zhang et al., 2014), with 20% of queries generating 80% of search volume, leaving 80% as long-tail queries. These long-tail queries are low in volume, transient, and rarely reappearing monthly. This transient nature limits the effectiveness of traditional exact match algorithms (Kharbanda et al., 2025; Mishra et al., 2025, 2024; Dahiya et al., 2021), as most labels won’t have future search volume. Due to the shifting dynamics of long-tail queries, predicting them individually as exact match keywords is nearly infeasible at time of inference. Instead, using other match types addresses this issue without explicit demand forecasting. Broad match keyphrases inherently address this challenge by capturing a wider set of related queries, thus effectively mitigating the unpredictability and transient nature of long-tail search patterns.
1.1.2. Management costs and Sales
@@
Managing sponsored search campaigns demands substantial effort from sellers, compounded by the extensive set of keyphrases requiring bids (Zhang et al., 2014; Amaldoss et al., 2016; Yang et al., 2021). Sellers prefer concise recommendations due to their perception and business demands (Mishra et al., 2025; Dey et al., 2025a). Other match type keyphrases are beneficial for targeting a wide array of exact queries, thereby reducing seller costs and involvement. Research suggests that when curation accuracy is high and seller bids remain competitive, broad match keyphrases yield higher sales (Yang et al., 2021).
1.2. Challenges for Other Match Types
1.2.1. Ground Truth Hurdles
@@
Several biases with click data have been well studied (Mishra et al., 2025; Chen et al., 2023; Vella, 1998; Gao et al., 2023; Dey et al., 2025b), these impact any form of supervised training for recommendation. Especially, non-exact match types are severely impacted by a sparsity of ground-truth labels. Unlike exact match keyphrases that are equivalent to buyer queries which can directly be recommended for the item by modeling on click data logs. Performance information of phrase/broad match types are only collected when the seller manually chooses them. Even though the sellers’ previous picks of successful broad match keyphrases could seem like a good basis for ground truth, they can be heavily affected by the seller set bid amounts and auctioning dynamics (Edelman et al., 2007).
@@
The limited availability of ground truth data leaves a significant portion of the feature space unmarked, requiring the incorporation of counterfactual data so models can learn significant patterns. Moreover, missing label biases (Steck, 2010) affects substantial counterfactual negative signals, as the absence of impressions or clicks doesn’t mean that the match type keyphrases are irrelevant. Consequently, the validity and dependability of historical performance metrics for assessing these recommendations become dubious. So, unlike exact match recommendations, which can be easily evaluated using classification metrics such as precision and recall score in a traditional setting (using ground truths), evaluating other match types present significant challenges.
1.2.2. Complexity of Reach
If a phrase/broad match keyphrase can be extended to a query using the definition in Section 1, then we term it as Reach. Theoretically, a keyphrase can reach a larger set of queries, assuming there is a finite number of actively buyer search queries. The broader the keyphrase the larger the reach. For example as shown in Figure 2, a generic broad type keyphrase ‘‘men’’ can reach thousands of queries pertaining to shoes, clothing, or accessories. Whereas a more specific ‘‘reebok men shoe’’ can reach only queries related to the specific brand, gender and accessory.
Although the broader keyphrase is beneficial to the seller by reaching a larger query set, it can lead to inefficient curation by flooding the matching platform illustrated in Figure 2. The matching for the more generic ‘‘mens’’ keyphrase will include a larger number of queries which are mostly irrelevant and increase computational costs for the platform. Also, due to the matching complexity of larger reach for the generic keyphrase, there is a high possibility of mismatching the seller’s item to an unprofitable buyer search query. This suboptimal targeting not only reduces the platform’s revenue, but also diminishes the seller’s return on ad spend (ROAS). Ideally, we want the broad match keyphrase to be closely matched with the reachable queries so that it takes less modifications to match the queries exactly.
1.2.3. Active Targeting
@@
Match type keyphrases recommended to sellers should be reachable to queries that are searched by buyers. Generated broad/phrase match keyphrases which are open vocabulary cannot reach such queries, though relevant to the item are considered futile leading to wastage of seller bids. [^1] See discussion of Out-of-Vocabulary models (Zhang et al., 2025; Simig et al., 2022; Gupta et al., 2021) and targeting in Graphex (Mishra et al., 2025). This is exacerbated by the transient queries mentioned in Section 1.1.1 that require keyphrases to target queries that buyers search for over time. Thus, there should a way of evaluating performance of broad matches on transient queries.
1.2.4. Execution Requirements
@@
Executing billions of item and query inferences within a few hours is paramount for the timely completion of daily tasks. Given the high memory demands, single-node data aggregation for inference poses a failure risk, while multi-node management elevates engineering expenses. Models need frequent updates for continuous integration of new data points, as discussed in Section 1.2.3. Deep learning networks like Large Language Models (LLMs) face scrutiny for high execution costs, notably due to their training and inference latency (Kaddour et al., 2023; Mishra et al., 2025; Sharma et al., 2023; Pope et al., 2023). Leading LLMs often rely on GPUs, raising acquisition costs and affecting both advertisers’ and platforms’ profit margins.
1.3. Scope and Contributions
In this work we only focus on Broad matches instead of Phrase matches  (eBay Inc., 2025; Inc., 2025) due to their wide applicability. We use our definition in Section 1 to streamline the evaluation; the actual implementation of the matching is platform-specific. We aim to tackle the challenges mentioned in Section 1.2, which we briefly describe here. We present a novel framework BroadGen for recommending broad match keyphrases for sellers and evaluating their performance. Due to the lack of ground truths per item (see Challenge 1.2.1), we incorporate relevant buyer queries as input (see Section 4.1) to determine the effective keyphrases for an item. This also helps us target active queries for Challenge 1.2.3. We also develop a unique evaluation methodology described in Sections 4.1 and  4.2 that works in absence of ground truths and assesses performance over time, emphasizing their adaptability and temporal resilience (Challenge 1.2.3).
@@ 1.3. Scope and Contributions
The Challenge 1.2.2 implies that a good broad match keyphrase should be neither too generic nor too specific, so that it can achieve a balance between reach and effectiveness. A “too specific” broad match keyphrase will behave like an exact match keyphrase which defeats the purpose. Our approach emphasizes avoiding overly broad suggestions, particularly in Section 3.4, with the aim of mitigating inefficiencies within both our systems and the business metrics of sellers. The metrics in Section 4.2 enable us to calculate the balance in effectiveness and efficiency of the recommended broad match keyphrases in relation to the search queries that it will be matched to. Our framework has been designed to be light and fast with minimal infrastructure costs and is easily deployable in a distributed manner (Challenge 1.2.4) to cater to millions of sellers with an inventory of 2.3 billion items daily.
2. Related Work
@@
Although broad match keyphrases are a crucial component, the research dedicated to specifically recommending novel broad match keyphrases is sparse. Traditionally, sellers have been responsible for selecting match types, with prior studies (Scholz et al., 2019) concentrating on creating keyphrases suitable for exact matches. A simplistic recommendation strategy has involved a rule-based Match Type Curation, which applies heuristics to assess whether an exact match type keyphrase, obtained through the current recall mechanism, has historically succeeded as a broad match keyphrase. This approach, however, faces intrinsic challenges: without historical data indicating that a seller has previously chosen the keyphrase as a broad match type — this method is unable to suggest the keyphrase due to a lack of pertinent performance indicators. Note that in this method the keyphrase is a query in itself.
Figure 3. Transfusion Algorithm Explained
@@
Some studies have focused on optimizing the matching of key-phrases to queries rather than effecting the sellers/advertiser’s selection of keyphrases. The BB-KSM model (Li and Yang, 2022) is a stochastic model that utilizes Markov Chain Monte Carlo (MCMC) to initially estimate keyphrase metrics such as impressions and CTR across match types, filling in incomplete data. Subsequently, it designates a match type to each keyphrase based on these metrics with the aim of maximizing a profit objective under budgetary constraints, employing a branch-and-bound algorithm for efficient decision space exploration. Similar optimization strategies have been documented for match types (Even Dar et al., 2009), which recommend using buyer search queries as broad match keyphrases rather than introducing new broad keyphrases. However, with the increasing variety of items and queries in e-commerce, this approach may be inefficient and expensive. The introduction of new queries and changing distributions (refer to Section 1.1.1) further reduces profitability. Studies like (Radlinski et al., 2008) apply query substitution methods (Jones and Fain, 2003; Jones et al., 2006) to connect broad keyphrases with buyer queries by preprocessing frequently searched queries and utilizing predefined substitution mappings.
@@
Transfusion is an eBay’s match type generative model developed based on its in-house NER (Named Entity Recognition) model (Xin et al., 2018; Palen-Michel et al., 2024) to recommend broad match type keyphrases especially for cold start items. The model generates keyphrases for a specified item utilizing the tokens present in its title. It ensures that these keyphrases align with popular and historically high-performing NER tag combinations within a particular meta category. For instance, as shown in Figure 3, the model identifies that the NER tag combination of product_name and storage_capacity is popular in this category based on the scoring function $f(X)$. The algorithm then construct 3 keyphrases: ‘‘iPhone 5S 16GB’’, ‘‘iPhone 5S 32GB’’, ‘‘iPhone 5S 64GB’’ based on the tokens corresponding to this NER tag combination from the item title. In this work, for comparison, we use a proxy for the rule-based match type curation and the Transfusion model against our model BroadGen.
3. Methodology
@@
Our framework requires a set of existing (buyer search) queries denoted as $Q$ and the item title $t$ for each item to recommend a set of keyphrases denoted as $K$. The goal is for the keyphrase set $K$ to be broad matched to the queries $Q$. Each query $q\in Q$, each keyphrase $k\in K$ and the title $t$ are in the form a list of tokens. [^2] A proprietary tokenization scheme handles normalization and stop word removal. We formulate the problem as String Clustering on the set of query strings ($Q$), to exploit the commonality within a cluster of queries and generate a representation for each cluster. The underlying intuition is that the cluster representation serves as the broad match keyphrase capable of addressing both the queries within the set and prospective queries beyond it. Our framework has three cores, Input generation (3.1) which generates the input matrix from the set of input queries. The Clustering core (3.2) which groups the queries using the input matrix. The Representation generation (3.3) constructs a representative ordered string for each cluster.
3.1. Input Generation

### v3

@@
Abstract
In the domain of sponsored search advertising, the focus of Keyphrase recommendation has largely been on exact match types, which pose issues such as high management expenses, limited targeting scope, and evolving search query patterns. Alternatives like Broad match types can alleviate certain drawbacks of exact matches but present challenges like poor targeting accuracy and minimal supervisory signals owing to limited advertiser usage. This research defines the criteria for an ideal broad match, emphasizing on both efficiency and effectiveness, ensuring that a significant portion of matched queries are relevant. We propose BroadGen, an innovative framework that recommends efficient and effective broad match keyphrases by utilizing historical search query data. Additionally, we demonstrate that BroadGen, through token correspondence modeling, maintains better query stability over time. BroadGen’s capabilities allow it to serve daily, millions of sellers at eBay with over 2.5 billion items.
I Introduction
Advertisers or sellers on e-commerce platforms are recommended keyphrases by the platform to increase their item visibility depending on buyer searches as shown in Figure 1(a). The sellers bid on relevant recommended keyphrases as well as their own custom keyphrases to increase their sales (Figure 1(b)). Keyphrase recommendation is the subject of a lot of research and well studied in several works [1, 2, 3, 4, 5, 6, 7], due to their challenging research and industrial importance. We term the recommendation to the seller as Keyphrase and the buyer search text as Query. Additionally, the platform matches the (recommended) keyphrases bid by the sellers to the buyer search queries [8].
@@ I Introduction
Figure 1: Screenshot of keyphrases for manual targeting in Promoted Listings Priority for eBay Advertising.
Each keyphrase is also tagged with a Match Type, which is used during the matching process; for this work we define them as:
  • Exact match type - The keyphrase is matched exactly to the query accommodating different cases, plurals, contractions, and similar substitutions handled by tokenization.
  • Phrase match type - It is more flexible by allowing modification to the keyphrase by in-position token expansion but maintaining the token order.
@@
For example, if a query is ‘‘reebok men’s shoes size 9’’, then the exact matching keyphrases can be ‘‘reebok men shoes size 9’’ or ‘‘Reebok men’s shoe size 9’’. The phrase match can be ‘‘reebok men shoe’’ while broad match can be ‘‘reebok shoe men’’. More examples of how different match types keyphrases for an item are matched to queries are shown in Figure 2. The platform can recommend a keyphrase with different match types to the sellers.
Previous research has focused on recommending exact match type keyphrases, popularly framing it as an Extreme Multi-Label Classification (XMC) [9, 10], where given an input item, they determine relevance to the labels which in this case are the buyer queries or keyphrases. Match types such as phrase and broad match keyphrases can alleviate problems associated with exact match type recommendations (see Section I-A). However, their recommendation strategies are under-researched as they are unlike exact match recommendations. Therefore, in this paper, we focus on other match type keyphrases by formulating the problem, designing solutions and providing insights learned in production.
Figure 2: Matching architecture
I-A Problems with Exact Match Type
I-A1 Shifting label distributions
@@
Analysis reveals that in sponsored search, queries adhere to the 80/20 rule where 20% of queries generate 80% of search volume leaving 80% as long-tail queries [11], which is prominent even in e-commerce platforms like eBay. These long-tail queries are low in volume, transient, and rarely reappearing monthly. This transient nature limits the effectiveness of traditional exact match algorithms [1, 2, 3, 4, 12], as most labels won’t have future search volume. Due to the shifting dynamics of long-tail queries, predicting them individually as exact match keywords is nearly infeasible at time of inference. Instead, using other match types addresses this issue without explicit demand forecasting. Broad match keyphrases inherently address this challenge by capturing a wider set of related queries, thus effectively mitigating the unpredictability and transient nature of long-tail search patterns.
I-A2 Management costs and Sales
@@
Managing sponsored search campaigns demands substantial effort from sellers, compounded by the extensive set of keyphrases requiring bids [11, 13, 14]. Sellers prefer concise recommendations due to their perception and business demands [2, 12]. Other match type keyphrases are beneficial for targeting a wide array of exact queries, thereby reducing seller costs and involvement. Research suggests that when curation accuracy is high and seller bids remain competitive, broad match keyphrases yield higher sales [14].
I-B Challenges for Other Match Types
Though non-exact match types are lucrative there are several challenges associated with recommending them.
I-B1 Ground Truth Hurdles
@@
Exact match keyphrases can be directly recommended by modeling on click data logs, since they correspond precisely to buyer queries. In contrast, phrase and broad match keyphrases are more difficult to model due to the sparsity and bias of available training signals. It suffers from sparsity as the performance information of phrase/broad match types are only collected when the seller manually chooses them. Although the sellers’ previous picks of successful broad match keyphrases could seem like a good basis for ground truth, they can be heavily affected by the seller set bid amounts and auctioning dynamics [15], thus making seller-based preference data a bad choice for modeling keyphrase recommendations [12].
@@
Several biases with click data have been well studied [2, 16, 17, 18, 8], these impact any form of supervised training for recommendation. The limited availability of ground truth data leaves a significant portion of the feature space unmarked, requiring the incorporation of counterfactual data so models can learn significant patterns. Moreover, missing label biases [19] affects substantial counterfactual negative signals, as the absence of impressions or clicks doesn’t mean that the match type keyphrases are irrelevant.
Consequently, the validity and dependability of historical performance metrics for assessing these recommendations become dubious. So, unlike exact match recommendations, which can be easily evaluated using classification metrics such as precision and recall score in a traditional setting (using ground truths), evaluating other match types presents significant challenges.
I-B2 Active Targeting
@@
Match type keyphrases recommended to sellers should be reachable to queries that are searched by buyers. Generated broad/phrase match keyphrases which are open vocabulary cannot reach such queries, though relevant to the item are considered futile leading to wastage of seller bids. [^1] See discussion of Out-of-Vocabulary models [20, 21, 22, 5] and targeting in Graphex [2]. This is exacerbated by the transient queries mentioned in Section I-A1 that require keyphrases to target queries that buyers search for over time. Thus, there should a way of evaluating performance of broad matches on transient queries.
I-B3 Complexity of Reach
If a phrase/broad match keyphrase can be extended to a query using the definition in Section I, then we term it as Reach. Theoretically, a keyphrase can reach a larger set of queries, assuming there is a finite number of actively buyer search queries. The broader the keyphrase the larger the reach. For example as shown in Figure 2, a generic broad type keyphrase ‘‘men’’ can reach thousands of queries pertaining to shoes, clothing, or accessories. Whereas a more specific ‘‘reebok men shoe’’ can reach only queries related to the specific brand, gender and accessory.
Although the broader keyphrase is beneficial to the seller by reaching a larger query set, it can lead to inefficient curation by flooding the matching platform illustrated in Figure 2. The matching for the more generic ‘‘men’’ keyphrase will include a larger number of queries which are mostly irrelevant and increase computational costs for the platform. Also, due to the matching complexity of larger reach for the generic keyphrase, there is a high possibility of mismatching the seller’s item to an unprofitable buyer search query. This suboptimal targeting not only reduces the platform’s revenue, but also diminishes the seller’s return on ad spend (ROAS). Ideally, we want the broad match keyphrase to be closely matched with the reachable queries so that it takes less modifications to match the queries exactly.
I-B4 Execution Requirements
@@
Executing billions of item and query inferences within a few hours is paramount for the timely completion of daily tasks. Given the high memory demands, single-node data aggregation for inference poses a failure risk, while multi-node management elevates engineering expenses. Models need frequent updates for continuous integration of new data points, as discussed in Section I-B2. Deep learning networks like Large Language Models (LLMs) face scrutiny for high execution costs, notably due to their training and inference latency [23, 2, 24, 25]. Leading LLMs often rely on GPUs, raising acquisition costs and affecting both advertisers’ and platforms’ profit margins.
I-C Scope and Contributions
@@
In this work we only focus on Broad matches instead of Phrase matches  [26, 27] due to their wide applicability. We use our definition in Section I to streamline the evaluation; the actual implementation of the matching is platform-specific. We aim to tackle the challenges mentioned in Section I-B, which we briefly describe here. We present a novel framework BroadGen for recommending broad match keyphrases to sellers and evaluate its performance. Due to the lack of ground truths per item (see Challenge I-B1), we incorporate relevant buyer queries as input (see Section IV-A) to determine the effective keyphrases for an item. This also helps us target active queries for Challenge I-B2. The Challenge I-B3 implies that a good broad match keyphrase should be neither too generic nor too specific, so that it can achieve a balance between reach and effectiveness. A “too specific” broad match keyphrase will behave like an exact match keyphrase which defeats the purpose. Our approach emphasizes avoiding overly broad suggestions, particularly in Section III-E, with the aim of mitigating inefficiencies within both our systems and the business metrics of sellers. Our framework has been designed to be light and fast with minimal infrastructure costs and is easily deployable in a distributed manner (Challenge I-B4) to cater to millions of sellers with an inventory of over 2.5 billion items daily. We also develop a unique evaluation methodology described in Sections IV-A and  IV-B that enables us to calculate the balance in effectiveness and efficiency of the recommended broad match keyphrases in relation to the search queries it will be matched to. It allows for absence of ground truths and assesses performance over time, emphasizing their adaptability and temporal resilience (Challenge I-B2).
II Related Work
@@
Although broad match keyphrases are a crucial component, the research dedicated to specifically recommending novel broad match keyphrases is sparse. Traditionally, sellers have been responsible for selecting match types, with prior studies [28] concentrating on creating keyphrases suitable for exact matches. A simplistic recommendation strategy has involved a rule-based Match Type Curation, which applies heuristics to assess whether an exact match type keyphrase, obtained through the current recall mechanism, has historically succeeded as a broad match keyphrase. This approach, however, faces intrinsic challenges: without historical data indicating that a seller has previously chosen the keyphrase as a broad match type — this method is unable to suggest the keyphrase due to a lack of pertinent performance indicators. Note that in this method the keyphrase is a query in itself.
Figure 3: Transfusion Algorithm Explained
@@
Some studies have focused on optimizing the matching of key-phrases to queries rather than effecting the sellers/advertiser’s selection of keyphrases. The BB-KSM model [29] is a stochastic model that utilizes Markov Chain Monte Carlo (MCMC) to initially estimate keyphrase metrics such as impressions and CTR across match types, filling in incomplete data. Subsequently, it designates a match type to each keyphrase based on these metrics with the aim of maximizing a profit objective under budgetary constraints, employing a branch-and-bound algorithm for efficient decision space exploration. Similar optimization strategies have been documented for match types [30], which recommend using buyer search queries as broad match keyphrases rather than introducing new broad keyphrases. However, with the increasing variety of items and queries in e-commerce, this approach may be inefficient and expensive. The introduction of new queries and changing distributions (refer to Section I-A1) further reduces profitability. Studies like [31] apply query substitution methods [32, 33] to connect broad keyphrases with buyer queries by preprocessing frequently searched queries and utilizing predefined substitution mappings.
@@
Transfusion is an eBay’s match type generative model developed based on its in-house NER (Named Entity Recognition) model [34, 35] to recommend broad match type keyphrases especially for cold start items. The model generates keyphrases for a specified item utilizing the tokens present in its title. It ensures that these keyphrases align with popular and historically high-performing NER tag combinations within a particular meta category. For instance, as shown in Figure 3, the model identifies that the NER tag combination of product_name and storage_capacity is popular in this category based on the scoring function $f(X)$. The algorithm then construct 3 keyphrases: ‘‘iPhone 5S 16GB’’, ‘‘iPhone 5S 32GB’’, ‘‘iPhone 5S 64GB’’ based on the tokens corresponding to this NER tag combination from the item title. In this work, for comparison, we use a proxy for the rule-based match type curation and the Transfusion model against our model BroadGen.
III Methodology
III-A Terms, Notations and Formulation
@@
Our framework requires a set of existing (buyer search) queries denoted as $Q$ and the item title $t$ for each item to recommend a set of keyphrases denoted as $K$. The goal is for the keyphrase set $K$ to be broad matched to the queries $Q$. Each query $q\in Q$, each keyphrase $k\in K$ and the title $t$ are in the form a list of tokens. [^2] A proprietary tokenization scheme handles normalization and stop word removal. We formulate the problem as String Clustering on the set of query strings ($Q$), to exploit the commonality within a cluster of queries and generate a representation for each cluster. The underlying intuition is that the cluster representation serves as the broad match keyphrase capable of addressing both the queries within the set and prospective queries beyond it. Our framework has three cores, Input generation (III-B) which generates the input matrix from the set of input queries. The Clustering core (III-C) which groups the queries using the input matrix. The Representation generation (III-D) constructs a representative ordered string for each cluster.
III-B Input Generation